OpenStudy (anonymous):

Give a description or schetch of∶ {z:Im(z ̅^2<1)} and {z ∶ (z-i) ^2 - (z+i)^2≤1 } as seperate sets of complex numbers and indicate as open or closed sets

OpenStudy (experimentx):

let z = x+iy z conjugate = x - iy im(z conjugate) = i 2xy = 2xy

OpenStudy (experimentx):

2xy <1 ... how does this look like? ... this is open set. I guess

OpenStudy (experimentx):

(z-i) ^2 - (z+i)^2 = -4ixy <= 1

OpenStudy (experimentx):

* -4iz <= 1

OpenStudy (experimentx):

probably this will give you a circle with radius 1/4 ... and this is closed set.

OpenStudy (experimentx):

*i guess

OpenStudy (anonymous):

I missed out on the two problems last night. (1-i)^(i*sqrt{2}) with i=imaginary number ; to prove that all values lie on a straight line and to find a branch of log(2z^2+1) that is analytic.

OpenStudy (experimentx):

that expression equals http://www.wolframalpha.com/input/?i=++%281-i%29^%28i*sqrt {2}%29+

OpenStudy (experimentx):

keep adding 2pi to it ... I don't think it will have any other root in 0, 2pi

OpenStudy (experimentx):

as for this problem .. log(2z^2+1) i think simplification would be the best thing you can do.

OpenStudy (experimentx):

\[ \log(Z) = \ln(z) + i arg(z+2 n\pi) \] probably use this formula to do it fase.

OpenStudy (experimentx):

it's quite awful to simplify it ... try following the yesterday's procedure.

OpenStudy (experimentx):

by letting z=x+iy << turn that expression into u(x,y) + i v(x,y)

OpenStudy (experimentx):

and about your earlier Q ... i think i kinda remembered few things http://openstudy.com/users/hennie#/updates/50430e6de4b000724d461e23

OpenStudy (experimentx):

after finding critical points, put those in place of coefficients ... and follow the usual procedure. find Eigen values ... once you know them, you know their nature.

OpenStudy (anonymous):

Thank you. This helped a lot.

OpenStudy (experimentx):

i hate non linear systems ...